A glass rod that has been charged to +13.0 NC touches a metal sphere. Afterward, the rod's charge is +7.0 NC.How many charged particles were transferred?

.
1
Expert's answer
2019-05-21T08:51:37-0400
"n=\frac{\Delta Q}{e}"

"n=\frac{13\cdot 10^{-9}-7\cdot 10^{-9}}{1.6\cdot 10^{-19}}=3.75\cdot 10^{10}"